OpenStudy (anonymous):

Which expression is equal to (2x + 5) · 8 using the commutative property? (5 + 2x) · 8 8 · 2x + 5 · 8 2x + (5 · 8) (2x + 5 · 8)

OpenStudy (kc_kennylau):

The commutative property means that a+b=b+a :)

OpenStudy (anonymous):

the commuative property is like 1+2=2+1

OpenStudy (anonymous):

The numbers on the left side are the same as the numbers on the right side.

OpenStudy (anonymous):

Is it A?

OpenStudy (kc_kennylau):

exactly :D